answersLogoWhite

0

With two dice the following totals are possible

....... Dice A .. Dice B

  1. 2 - 1 ..... 1
  2. 3 - 1 ..... 2
  3. 3 - 2 ..... 1
  4. 4 - 1 ..... 3
  5. 4 - 2 ..... 2
  6. 4 - 3 ..... 1
  7. 5 - 1 ..... 4
  8. 5 - 2 ..... 3
  9. 5 - 3 ..... 2
  10. 5 - 4 ..... 1
  11. 6 - 1 ..... 5
  12. 6 - 2 ..... 4
  13. 6 - 3 ..... 3
  14. 6 - 4 ..... 2
  15. 6 - 5 ..... 1
  16. 7 - 1 ......6
  17. 7 - 2 ..... 5
  18. 7 - 3 ..... 4
  19. 7 - 4 ..... 3
  20. 7 - 5 ..... 2
  21. 7 - 6 ..... 1
  22. 8 - 2 ......6
  23. 8 - 3 ..... 5
  24. 8 - 4 ..... 4
  25. 8 - 5 ..... 3
  26. 8 - 6 ..... 2
  27. 9 - 3 ..... 6
  28. 9 - 4 ..... 5
  29. 9 - 5 ..... 4
  30. 9 - 6 ..... 3
  31. 10 - 4 ..... 6
  32. 10 - 5 ..... 5
  33. 10 - 6 ..... 4
  34. 11 - 5 ..... 6
  35. 11 - 6 ..... 5
  36. 12 - 6 ..... 6

So, there are 36 different combinations of how the dice might land and 6 of those total 7. Therefore it is 6 chances in 36 that you'll get 7, which reduced to 1 in 6.

User Avatar

Wiki User

∙ 15y ago

What else can I help you with?

Related Questions

When two dice are rolled what is the probability of getting a sum of 14?

You can't get 14 with two regular six-sided dice ! The highest you can get with one throw is 12.


If 11 is the sum and you roll two dice what is the probability of getting 11 if you toss the dice 54 times?

The probability of getting 11 with one throw of 2 dice is 1/6*1/6*2 = 1/18 So the probability of not getting 11 with 1 throw of the dice is 17/18. Tossing the dice 54 times, the probability of not getting 11 54 times is (17/18)54 = 0.0456... So the probability of at least 1 roll of 11 is 1 - 0.0456 = 0.954


What is the probability that the sum is not seven if you throw 2 dice?

It is 30/36 = 5/6


What are the chances of getting a sum of 70 on two dice?

The probability is zero. The largest possible sum on two dice is 12.


When two dice are rolled find the probability of getting a sum of greater than 15?

Two standard dice cannot have a sum greater than 12. The probability, then, of getting a sum greater than 15 is zero.


When two dice are rolled find the probability of getting a sum less than 13?

Assuming these are regular dice, the probability is 1.


What is the probability of obtaining a sum of 11 in a single throw with 3 dice?

The probability is 27/216 = 3/8 = 0.125


What is the probability of getting the sum of 8 from a pair of dice?

It is 5/36.


When two dice are rolled what is the probability of not getting a sum of 12?

The probability of rolling 12 with 2 dice is 1 in 36. The probability of not rolling 12 with 2 dice is 35 in 36.


When two dice are rolled the probability of getting a sum of 16?

With two normal six-sided dice, the maximum sum can only be 12 (6 x 2 dice), so there is zero probability of rolling a sum that adds to 16.


How many times must you throw two standard dice to be sure you get the same sum at least twice?

The probability of gettting a particular sum on a standard set of dice depends on which sum you are seeking. For example, the sums of 2 and 12 have a probability of 1 in 36, or about 0.0278; while the sum of 7 has a probability of 6 in 36, or 1 in 6, or about 0.167.Specifically answering the question; it is not possible to guarantee a particular outcome in a random throw, or in a series of random throws, of the dice. You can only talk about probability. Let's take the worst case of trying to throw a 2 or a 12. Even of you throw the dice 100 times, the probability is only 0.0278100, or about 2.34 x 10-156 that you will not throw the 2 or 12; so, the probability is extremely good that you will throw a 2 or 12 in 100 throws, but it is not guaranteed. That's the thing about probability.


When two dice are rolled find the probability of getting a 6 or 7?

Probability that the sum is 6 = 5/36 Probability that the sum is 7 = 6/36

Trending Questions
A family has 8 girls and 4 boys. A total of 2 children must be chosen to speak on the behalf of the family at a local benefit. What is the probability that 1 girl and 1 boy will be chosen? What is the value of p if 9p 162? Does a greater tolerance for alcohol lower a persons BAC? What does XVVMMIX mean? What is the volume of a six sided dice? What month had the highest numberEach year an average of eight teens 16 to 19 were killed every day from motor vehicle crashes. of teenage crash deaths of any month.? What is the probability of drawing a black card and a 7? IQ test scores are normally distributed with a mean of 100 and a standard deviation of 15 An individual's IQ score is found to be 110 Find the z-score corresponding to this value? How many pounds are in 500 metric tons? What does one twelfth look like on a pie chart? The probability density function has what value in the interval between 20 and 28? What is the importance of averages? If the probability that an event will happen is .25 or 25 percent what is the probability that it will not happen? What is a diagram that shows relationships among sets of objects? What does trend line represent in a forex chart? When using the distribution of sample mean to estimate the population mean what is the benefit of using larger sample sizes? What is an IQ of 165 mean? How many times does it take for you to get an STD and the medicine doesnt work? If you take Statistics AP in high school do you have to take it again in college? What is the ample weight for someone 5 foot eleven inches?